OpenStudy (anonymous):

if i am given a function w(x,y,z) as well as three other equations x=..., y=..., z=... and asked to find dw/dy how to i go about doing this

OpenStudy (anonymous):

w=x^2+y^2+z^2, x=uv, y=ucos(v), z=usin(v) find dw/dy when (u,v)=(1,0)